# The Intelligible and Effective Graph Neural Additive Networks

### v3

@@ 3 Graph Neural Additive Networks
In this section, we introduce the Graph Neural Additive Networks (GNAN). We begin by defining some essential notation. A graph $G$ has a set of $N$ vertices, where each vertex is associated with a $d$-dimensional feature vector. Specifically, $\mathbf{x}_{i}\in\mathbb{R}^{d}$ represents the feature vector of the $i$’th node in $G$. We define the distance $\text{dist}\left(j,i\right)$ between node $j$ and node $i$ within the graph $G$ as the number of edges in the shortest path from $j$ to $i$. This definition implies that the distance from a node to itself is zero. In cases where no path exists from $j$ to $i$, we set $\text{dist}\left(j,i\right)=\infty$. For enhanced readability, we denote vectors in boldface, and an entry $k$ of a vector $\mathbf{x}$ is denoted by $[\mathbf{x}]_{k}$. We begin by describing GNAN for applications such as binary classification and regression where the model output is one-dimensional. At the end of this section, we discuss extensions to scenarios such as multi-class classification, where the model output is multi-dimensional.
@@
$[\mathbf{h}_{i}]_{k}=\sum_{j=1}^{N}\frac{1}{\#\text{dist}_{i}(j,i)}\cdot\rho\left(\frac{1}{1+\text{dist}\left(j,i\right)}\right)\cdot f_{k}\left(\left[\mathbf{x}_{j}\right]_{k}\right)$
where #$\#\text{dist}_{i}(j,i)$ represents the number of nodes at distance $\text{dist}\left(j,i\right)$ from node $i$. The underlying rationale for this definition is as follows: each node’s $k$’th feature is transformed by a shape function $f_{k}$, independently from other features. The effect the $k$’th feature value of node $j$ has on node $i$’s representation is influenced by their distance. Specifically, if $\text{dist}\left(j,i\right)=l$, then the cumulative influence of all nodes at distance $l$ from node $i$ is captured by $\rho(\nicefrac{{1}}{{(1+l)}})$. This is achieved by the normalization term $\nicefrac{{1}}{{\#\text{dist}_{i}(j,i)}}$. Here, $\rho$’s argument $\nicefrac{{1}}{{(1+l)}}$ scales the distance such that a distance of 00 (the self-distance of a node) is mapped to $1$, and an infinite distance, which implies no path exists, is scaled to 00. Thus, $\rho$ spans the interval $[0,1]$.
@@
Here $[\mathbf{h}_{i}]_{k}$ contains the influence of the $k$’th feature in node $i$ on the prediction. However, from the definition of $[\mathbf{h}_{i}]_{k}$ we see that it serves as a mediator for influences of all other nodes. Therefore, the influence of node $j$ on feature $k$ in the final graph representation can be obtained by:
@@
$[\mathbf{h}_{i}]_{k}=\sum_{j=1}^{N}\frac{1}{\#\text{dist}_{i}(j,i)}\cdot\rho\left(\frac{1}{1+\text{dist}\left(j,i\right)}\right)\odot f_{k}([\mathbf{x}_{j}]_{k})~{}~{}~{}.$
@@ 4 Inteligibility
Our initial examples focus on the task of detecting mutation-causing molecules using the Mutagenicity dataset [47]. In this task, molecules are modeled as graphs where nodes correspond to atoms and edges to connections between these atoms. Each atom type is represented by a 14-dimensional one-hot encoding. A GNAN model trained on this dataset is illustrated in Figure 1. On the left, the function $\rho$ is presented, demonstrating how distance impacts prediction, with a clear diminishing influence of more distant atoms. On the right, the shape functions for the features are displayed. Given that the features are binary, each shape function manifests only two values: one when the feature is 00 (indicating that the atom is not of the specified type), and another when the feature is $1$ (indicating that the atom is of the specified type). Defining $b=\sum_{k}f_{k}(0)$ as the bias term allows us to set $f_{k}(0)=0$ for each $k$, thereby enabling the plotting of only $f_{k}(1)$. The graphical representation reveals that atoms such as Ca, Na, and Li are predicted to correlate with an increased mutagenicity effect, whereas N and P atoms are predicted to be associated with a slight protective effect.
It is essential to emphasize that Figure 1 displays the entire model comprehensively. This means that combined with the value of the bias term, which is $-5.6672$ in this case, every crucial detail needed to understand and utilize this model for predictions is contained within this single figure. This stands in stark contrast to methods like feature importance, which offer a limited perspective on models. While the figure provides complete information about the model, presenting additional views can sometimes be helpful.
@@ 4.1 Local Explanations
Figure LABEL:fig:local presents two such examples. In Figure LABEL:fig:carbon_ring the carbon (red) atoms play the most significant role, and a carbon ring (red cycle) is highlighted. In Figure LABEL:fig:no2, a group of $NO_{2}$ (grey and green subgraph) is shown to be relatively important for predicting the molecule as mutagenic. Both carbon rings and $NO_{2}$ groups are well-known for their mutagenic effects [51], making them frequently discussed in the literature on explainable GNNs [39, 41].
@@ 5 Empirical evaluation
Thr $\mu$ ,$\alpha$ ,$\alpha_{HOMO}$ [60] datasets are long-range molecular property prediction regression tasks, over the large-scale QM9 molecular dataset.
@@ D.3 Hyperparameters
All GNNs (excluded GNAN) use ReLU activations with $\{3,5\}$ layers and $64$ hidden channels. They were trained with Adam optimizer over $1000$ epochs and early on the validation loss with a patient of $100$ steps, eight Decay of $41e-$ 4, learning rate in $\{1e-3,1e-4$}, dropout rate in $\{0,0.5\}$, and a train batch size of $32$.
In GNAN, all the feature and distance networks use ReLU activations with $\{3,5\}$ layers and $\{64,32\}$ hidden channels. They were trained with Adam optimizer over $1000$ epochs weight decay of $40,5e-$ 4, learning rate in $\{1e-2,1e-3$}, dropout rate in $\{0,0.6\}$.

### v4

@@ 3 Graph Neural Additive Networks
In this section, we introduce the Graph Neural Additive Networks (GNAN). We begin by defining some essential notation. A graph $G$ has a set of $N$ vertices, where each vertex is associated with a $d$-dimensional feature vector. Specifically, $\mathbf{x}_{i}\in\mathbb{R}^{d}$ represents the feature vector of the $i$’th node in $G$. We define the distance $\text{dist}\left(j,i\right)$ between node $j$ and node $i$ within the graph $G$ as the number of edges in the shortest path from $j$ to $i$. This definition implies that the distance from a node to itself is zero. In cases where no path exists from $j$ to $i$, we set $\text{dist}\left(j,i\right)=\infty$. For enhanced readability, we denote vectors in boldface, and an entry $k$ of a vector $\mathbf{x}$ is denoted by $\mathbf{x}]_{k}$. We begin by describing GNAN for applications such as binary classification and regression where the model output is one-dimensional. At the end of this section, we discuss extensions to scenarios such as multi-class classification, where the model output is multi-dimensional.
@@
$\mathbf{h}_{i}]_{k}=\sum_{j=1}^{N}\frac{1}{\#\text{dist}_{i}(j,i)}\cdot\rho\left(\frac{1}{1+\text{dist}\left(j,i\right)}\right)\cdot f_{k}\left(\left[\mathbf{x}_{j}\right]_{k}\right)$
where $\text{dist}_{i}(j,i)$ represents the number of nodes at distance $\text{dist}\left(j,i\right)$ from node $i$. The underlying rationale for this definition is as follows: each node’s $k$’th feature is transformed by a shape function $f_{k}$, independently from other features. The effect the $k$’th feature value of node $j$ has on node $i$’s representation is influenced by their distance. Specifically, if $\text{dist}\left(j,i\right)=l$, then the cumulative influence of all nodes at distance $l$ from node $i$ is captured by $\rho(\nicefrac{{1}}{{(1+l)}})$. This is achieved by the normalization term $\nicefrac{{1}}{{\#\text{dist}_{i}(j,i)}}$. Here, $\rho$’s argument $\nicefrac{{1}}{{(1+l)}}$ scales the distance such that a distance of 0 (the self-distance of a node) is mapped to $1$, and an infinite distance, which implies no path exists, is scaled to 0. Thus, $\rho$ spans the interval $[0,1]$.
@@
Here $\mathbf{h}_{i}]_{k}$ contains the influence of the $k$’th feature in node $i$ on the prediction. However, from the definition of $\mathbf{h}_{i}]_{k}$ we see that it serves as a mediator for influences of all other nodes. Therefore, the influence of node $j$ on feature $k$ in the final graph representation can be obtained by:
@@
$\mathbf{h}_{i}]_{k}=\sum_{j=1}^{N}\frac{1}{\#\text{dist}_{i}(j,i)}\cdot\rho\left(\frac{1}{1+\text{dist}\left(j,i\right)}\right)\odot f_{k}([\mathbf{x}_{j}]_{k})~~~.$
@@ 4 Inteligibility
Our initial examples focus on the task of detecting mutation-causing molecules using the Mutagenicity dataset [47]. In this task, molecules are modeled as graphs where nodes correspond to atoms and edges to connections between these atoms. Each atom type is represented by a 14-dimensional one-hot encoding. A GNAN model trained on this dataset is illustrated in Figure 1. On the left, the function $\rho$ is presented, demonstrating how distance impacts prediction, with a clear diminishing influence of more distant atoms. On the right, the shape functions for the features are displayed. Given that the features are binary, each shape function manifests only two values: one when the feature is 0 (indicating that the atom is not of the specified type), and another when the feature is $1$ (indicating that the atom is of the specified type). Defining $b=\sum_{k}f_{k}(0)$ as the bias term allows us to set $f_{k}(0)=0$ for each $k$, thereby enabling the plotting of only $f_{k}(1)$. The graphical representation reveals that atoms such as Ca, Na, and Li are predicted to correlate with an increased mutagenicity effect, whereas N and P atoms are predicted to be associated with a slight protective effect.
It is essential to emphasize that Figure 1 displays the entire model comprehensively. This means that combined with the value of the bias term, which is 5.6672-5.6672 in this case, every crucial detail needed to understand and utilize this model for predictions is contained within this single figure. This stands in stark contrast to methods like feature importance, which offer a limited perspective on models. While the figure provides complete information about the model, presenting additional views can sometimes be helpful.
@@ 4.1 Local Explanations
Figure 7(a) presents two such examples. In Figure LABEL:fig:carbon_ring the carbon (red) atoms play the most significant role, and a carbon ring (red cycle) is highlighted. In Figure LABEL:fig:no2, a group of $NO_{2}$ (grey and green subgraph) is shown to be relatively important for predicting the molecule as mutagenic. Both carbon rings and $NO_{2}$ groups are well-known for their mutagenic effects [51], making them frequently discussed in the literature on explainable GNNs [39, 41].
(a) Local explanations of two molecules from the Mutagenicity datasets, through visualizations of the molecules. The area of each atom corresponds to the node importance according to Equation 3.
@@ 5 Empirical evaluation
Thr $\mu$ $\alpha$ $\alpha_{HOMO}$ [60] datasets are long-range molecular property prediction regression tasks, over the large-scale QM9 molecular dataset.
@@ D.3 Hyperparameters
All GNNs (excluded GNAN) use ReLU activations with $\{3,5\}$ layers and $64$ hidden channels. They were trained with Adam optimizer over $1000$ epochs and early on the validation loss with a patient of $100$ steps, eight Decay of 1e41e-4, learning rate in $\{1e-3,1e-4$}, dropout rate in $\{0,0.5\}$, and a train batch size of $32$.
In GNAN, all the feature and distance networks use ReLU activations with $\{3,5\}$ layers and $\{64,32\}$ hidden channels. They were trained with Adam optimizer over $1000$ epochs weight decay of 0,5e40,5e-4, learning rate in $\{1e-2,1e-3$}, dropout rate in $\{0,0.6\}$.